# Second- and third-order properties of multidimensional Langevin equations

### v16

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
where the time-derivative is interpreted in the Itô sense, and [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] is the covariation defined by:
(6) $[x,y](t)=\lim_{\Delta\to 0}\sum_{n=0}^{N-1}(x(t_{n+1})-x(t_{n}))(y(t_{n+1})-y(t_{n}))$
@@
($\mathsf{H}$ denoting Hermitian conjugate and $\Re$ denoting the real part), therefore $-\Re\lambda$ is bounded by the smallest and largest eigenvalues of $\mathbf{D}$ [14]. In addition, the trace of $\mathbf{D}\mathbf{C}^{-1}$ equals negative the trace of $\mathbf{A}$.
@@ 2.2. Comparison of experimental data with theoretical predictions
The second-rank mixed tensor $\mathbf{M}\mathbf{C}^{-1}$ has real eigenvalues because it is symmetric in “covariance-identity” coordinates. We see that the r.h.s. of the above is the sum of the squared eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$. We may use this quantity to evaluate quantitative significance [^5] It may be noticed that in “covariance-identity” coordinates, the “cross-covariances” ($M^{ij}$ where $i\neq j$) are counted with double the contribution to $\operatorname{tr}(\bm{\Gamma})$ as compared to “self-covariances” ($M^{ij}$ where $i=j$). This is necessary for a coordinate-independent evaluation. We may illustrate this on a two-dimensional example, with variables $x$ and $y$. We fix $\langle x^{2}\rangle=\langle y^{2}\rangle=1$, while $\langle xy\rangle$ is allowed to deviate from 0. Changing coordinates to $x^{\prime}:=(x+y)/\sqrt{2}$ and $y^{\prime}=(y-x)/\sqrt{2}$, we obtain $\langle{x^{\prime}}^{2}\rangle=1+\langle xy\rangle$, $\langle{y^{\prime}}^{2}\rangle=1-\langle xy\rangle$, and $\langle x^{\prime}y^{\prime}\rangle=0$. We thus see that if squared deviations of $\langle{x^{\prime}}^{2}\rangle$ and $\langle{y^{\prime}}^{2}\rangle$ to 1 are to be added, we must add the contribution due to $\langle xy\rangle$ twice. Furthermore, for a two-dimensional Ornstein–Uhlenbeck process with $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$, the time-averages for a trajectory of finite length, denoted by overlines, satisfy $\langle(\overline{x^{2}}-1)^{2}\rangle=2\langle(\overline{xy})^{2}\rangle$..
First, however, we must address a problem with Eq. (12), which is that it does not obey symmetry in the indices. We may remedy this by symmetrizing its r.h.s. To justify this choice, consider a complete set of left eigenvectors of $\bm{\Gamma}$. Such a set exists because a rank-one matrix can always be diagonalized. In dimension $d$, there are $d(d-1)/2$ linearly independent left eigenvectors such that $v_{ji}=-v_{ij}$; we call these “antisymmetric” left eigenvectors and are associated with a zero eigenvalue. We may assume that they are included our complete set of left eigenvectors. Let $v_{ij}^{(k)}$ be the remaining left eigenvectors of $\bm{\Gamma}$ with eigenvalues $\mu^{(k)}$ ($k$ taking $d(d+1)/2$ values), so that:
@@ 2.3.1. Stochastic rotation frequencies
Now, we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\bm{\Omega}$ has a single pair of purely imaginary eigenvalues ±$\pm i\omega_{\mathrm{stoch}}$. We will see that the possible values of $\omega_{\mathrm{stoch}}$ are restricted by the eigenvalues of $\mathbf{A}$. The angular momentum is $L:=\langle x\dot{y}-\dot{x}y\rangle$. It is related to the matrix $\bm{\Omega}$ by the relation:
@@ 2.3.2. Quantitative significance of broken detailed balance
where the last equality follows from the Lyapunov equation. (We may understand the factor of 2 in the definition of $\mathbf{H}$ by considering that the symmetric part of $\mathbf{A}\mathbf{C}$ is $-\mathbf{D}$ by the Lyapunov equation, so the antisymmetric part should be compared to $\mathbf{D}$.) The entropy production rate $\dot{S}$ is related to $\mathbf{H}$ as $\dot{S}=-\operatorname{tr}(\mathbf{A}\mathbf{H})$ [13]. By a similar argument as in [13], in coordinates in which $\mathbf{D}=\mathds{1}$, we see that $\mathbf{H}$ is antisymmetric and therefore, like $\bm{\Omega}$, has purely imaginary complex conjugate eigenvalues. As an alternative derivation, following [3], we may consider that $\mathbf{D}^{-1/2}\mathbf{H}\mathbf{D}^{1/2}$ is antisymmetric and therefore $\mathbf{H}$ is similar to an antisymmetric matrix. In two dimensions the (dimensionless) eigenvalues ±$\pm ih$ are given by a simple formula similar to that of $\omega_{\mathrm{stoch}}$:
@@
If $h$ is at least comparable to unity (more precisely, $1/\sqrt{2}$; see end of subsubsection), then we consider that broken detailed balance is significant. We may illustrate this criterion in the cases of real or complex eigenvalues of $\mathbf{A}$. For two real eigenvalues $-\lambda_{x}$ and $-\lambda_{y}$ of $\mathbf{A}$ (Eq. (32)),
@@
For the case of complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ with $D^{xy}=0$ (Eq. (34)),
@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) in [9] to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. To explore the impact of dimension, we explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^8] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9]. Also, we have neglected the inference of the true mean $\langle\mathbf{x}\rangle$ and have assumed this as known to be $\mathbf{0}$. This simplifies the calculations while retaining interesting effects.:
@@
The $n=0$ term contributes $-\mathds{1}$ to $\langle\widehat{\mathbf{A}}(T)\rangle$. For the contribution of the $n=1$ term, we have from Gaussianity and Isserlis’s theorem (also known as Wick’s theorem [16]):
@@
.$\Big{.}{\qquad{}+\langle\dot{x}^{i}(\tau)x^{l^{\prime}}(\tau^{\prime\prime})\rangle(\langle x^{j}(\tau)x^{k}(\tau^{\prime})\rangle\langle x^{k^{\prime}}(\tau^{\prime\prime})x^{l}(\tau^{\prime})\rangle+\langle x^{j}(\tau)x^{l}(\tau^{\prime})\rangle\langle x^{k}(\tau^{\prime})x^{k^{\prime}}(\tau^{\prime\prime})\rangle)}\Big{]}.$
@@
(52) $-\delta^{ij}(\delta^{kk^{\prime}}\delta^{ll^{\prime}}+\delta^{kl^{\prime}}\delta^{lk^{\prime}})\frac{2T+e^{-2T}-1}{2T^{2}}$
@@
Figure 2. Effect of dimensionality on mean of (inferred) $\widehat{A}^{i}_{i}$ ($i$ not summed) when $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$.
@@
.$\Big{.}{\qquad{}+\langle\dot{x}^{i}(\tau)x^{l}(\upsilon)\rangle(\langle x^{j}(\tau)\dot{x}^{i^{\prime}}(\tau^{\prime})\rangle\langle x^{j^{\prime}}(\tau^{\prime})x^{k}(\upsilon)\rangle+\langle x^{j}(\tau)x^{j^{\prime}}(\tau^{\prime})\rangle\langle\dot{x}^{i^{\prime}}(\tau^{\prime})x^{k}(\upsilon)\rangle)}\Big{]}.$
@@
×$\qquad{}\times(\langle x^{k}(\upsilon)x^{k^{\prime}}(\upsilon^{\prime})\rangle\langle x^{l}(\upsilon)x^{l^{\prime}}(\upsilon^{\prime})\rangle+\langle x^{k}(\upsilon)x^{l^{\prime}}(\upsilon^{\prime})\rangle\langle x^{l}(\upsilon)x^{k^{\prime}}(\upsilon^{\prime})\rangle)+\mathcal{O}\left({\frac{1}{T^{2}}}\right)$
@@
Figure 3. Effect of dimensionality on variance of (inferred) $\widehat{A}^{i}_{i}$ ($i$ not summed) when $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$.
@@
Figure 4. Effect of dimensionality on variance of inferred drift ($\widehat{A}^{i}_{k}\widehat{A}^{i}_{l}\widehat{C}^{kl}$, $i$ not summed) when $-\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$.
@@ 3.1. Preliminaries: Modeling, covariance functions, and detailed balance
 [^10] As with $\mathbf{x}$, we have omitted a possible additive constant in $\dot{y}$. A system with an additive constant may be transformed into Eq. (60) by considering instead the quantity $y(t)-ct$, where $c$ is a constant (in this case, $c=\langle\dot{y}\rangle$). This simplification is again used throughout. Note that such a constant must vanish when detailed balance is satisfied, since we must have in that case $\langle\dot{y}\rangle=0$.
@@ 3.3. Comparison of experimental data with theoretical predictions
where $\Delta_{\textrm{exp\textendash theo}}$ denotes the deviation of experimental to theoretical values. As mentioned for the angular momentum, if the time-step is not small, then on the r.h.s. of the above, it may be suitable to use discrete-time estimators of the diffusion. In case the angular momenta exceed the “reference” values (Eq. (43)) by some factor, we may need to relax the above comparison by the same factor, not just for the antisymmetric part but also for the symmetric part. We may illustrate this by way of two-dimensional linear Gaussian dynamics with complex conjugate eigenvalues $-\lambda\pm i\omega$ of $\mathbf{A}$ (Eq. (34)) and $D^{xx}=D^{yy}$, in which case:
@@
[…]
(84) $2D^{yz}=\lim_{\tau\to 0^{+}}\frac{\langle(y(\tau)-y(0))(z(\tau)-z(0))\rangle}{\tau}=\text{``$\langle-\dot{y}z-y\dot{z}\rangle$''}.$ ” .
Since there is no angular momentum between $y$ and $z$, we assign $\langle\dot{y}z\rangle$ and $\langle y\dot{z}\rangle$ the same (fictional) value, $-D^{yz}$. For non-zero time lags, we integrate:
[…]
The above is equal to $-\langle(z(\tau)-z(0))\circ\dot{y}(\tau)\rangle$. From the velocity covariance function $\langle\dot{y}(t)\dot{z}(0)\rangle$ (where $t$ is allowed to run negative), we can compute the expected value of any second-order quantity, such as the two-variable analogue of the mean squared displacement:
@@
where we re-indexed $m=n+k$. Thus, although the estimate Eq. (90) for the velocity autocovariance function is self-averaging for any given finite time lag, its sum over all time lags is not [^11] In the case where $\langle\Delta x\rangle$ is treated as known and can be used in place of $\overline{\Delta x}$ in Eq. (92), the resulting estimate of long-time diffusivity is $(N/2\Delta t)(\overline{\Delta x}-\langle\Delta x\rangle)^{2}$, which has the correct expectation value, but is not self-averaging.. If we denote the limits of summation by ±$\pm M$, then the limits $M\to\infty$, $N\to\infty$ do not commute. The same considerations apply when integrating the covariance functions of stationary variables.

### v17

@@ 2.1. Preliminaries: Definition, covariance functions, and detailed balance
where the time-derivative is interpreted in the Itô sense, and $\cdot$$\cdot$] is the covariation defined by:
(6) $x,y](t)=\lim_{\Delta\to 0}\sum_{n=0}^{N-1}(x(t_{n+1})-x(t_{n}))(y(t_{n+1})-y(t_{n}))$
@@
($\mathsf{H}$ denoting Hermitian conjugate and $\Re$ denoting the real part), therefore $\Re\lambda$ is bounded by the smallest and largest eigenvalues of $\mathbf{D}$ [14]. In addition, the trace of $\mathbf{D}\mathbf{C}^{-1}$ equals negative the trace of $\mathbf{A}$.
@@ 2.2. Comparison of experimental data with theoretical predictions
The second-rank mixed tensor $\mathbf{M}\mathbf{C}^{-1}$ has real eigenvalues because it is symmetric in “covariance-identity” coordinates. We see that the r.h.s. of the above is the sum of the squared eigenvalues of $\mathbf{M}\mathbf{C}^{-1}$. We may use this quantity to evaluate quantitative significance [^5] It may be noticed that in “covariance-identity” coordinates, the “cross-covariances” ($M^{ij}$ where $i\neq j$) are counted with double the contribution to $\operatorname{tr}(\bm{\Gamma})$ as compared to “self-covariances” ($M^{ij}$ where $i=j$). This is necessary for a coordinate-independent evaluation. We may illustrate this on a two-dimensional example, with variables $x$ and $y$. We fix $\langle x^{2}\rangle=\langle y^{2}\rangle=1$, while $\langle xy\rangle$ is allowed to deviate from 0. Changing coordinates to $x^{\prime}:=(x+y)/\sqrt{2}$ and $y^{\prime}=(y-x)/\sqrt{2}$, we obtain $\langle{x^{\prime}}^{2}\rangle=1+\langle xy\rangle$, $\langle{y^{\prime}}^{2}\rangle=1-\langle xy\rangle$, and $\langle x^{\prime}y^{\prime}\rangle=0$. We thus see that if squared deviations of $\langle{x^{\prime}}^{2}\rangle$ and $\langle{y^{\prime}}^{2}\rangle$ to 1 are to be added, we must add the contribution due to $\langle xy\rangle$ twice. Furthermore, for a two-dimensional Ornstein–Uhlenbeck process with $\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$, the time-averages for a trajectory of finite length, denoted by overlines, satisfy $\langle(\overline{x^{2}}-1)^{2}\rangle=2\langle(\overline{xy})^{2}\rangle$..
First, however, we must address a problem with Eq. (12), which is that it does not obey symmetry in the indices. We may remedy this by symmetrizing its r.h.s. To justify this choice, consider a complete set of left eigenvectors of $\bm{\Gamma}$. Such a set exists because a rank-one matrix can always be diagonalized. In dimension $d$, there are d(d1)/2d(d-1)/2 linearly independent left eigenvectors such that $v_{ji}=-v_{ij}$; we call these “antisymmetric” left eigenvectors and are associated with a zero eigenvalue. We may assume that they are included our complete set of left eigenvectors. Let $v_{ij}^{(k)}$ be the remaining left eigenvectors of $\bm{\Gamma}$ with eigenvalues $\mu^{(k)}$ ($k$ taking $d(d+1)/2$ values), so that:
@@ 2.3.1. Stochastic rotation frequencies
Now, we specialize to the two-dimensional case, where $\mathbf{x}=(x,y)^{\mathsf{T}}$. Here, $\bm{\Omega}$ has a single pair of purely imaginary eigenvalues $\pm i\omega_{\mathrm{stoch}}$. We will see that the possible values of $\omega_{\mathrm{stoch}}$ are restricted by the eigenvalues of $\mathbf{A}$. The angular momentum is $L:=\langle x\dot{y}-\dot{x}y\rangle$. It is related to the matrix $\bm{\Omega}$ by the relation:
@@ 2.3.2. Quantitative significance of broken detailed balance
where the last equality follows from the Lyapunov equation. (We may understand the factor of 2 in the definition of $\mathbf{H}$ by considering that the symmetric part of $\mathbf{A}\mathbf{C}$ is $\mathbf{D}$ by the Lyapunov equation, so the antisymmetric part should be compared to $\mathbf{D}$.) The entropy production rate $\dot{S}$ is related to $\mathbf{H}$ as $\dot{S}=-\operatorname{tr}(\mathbf{A}\mathbf{H})$ [13]. By a similar argument as in [13], in coordinates in which $\mathbf{D}=\mathds{1}$, we see that $\mathbf{H}$ is antisymmetric and therefore, like $\bm{\Omega}$, has purely imaginary complex conjugate eigenvalues. As an alternative derivation, following [3], we may consider that $\mathbf{D}^{-1/2}\mathbf{H}\mathbf{D}^{1/2}$ is antisymmetric and therefore $\mathbf{H}$ is similar to an antisymmetric matrix. In two dimensions the (dimensionless) eigenvalues $\pm ih$ are given by a simple formula similar to that of $\omega_{\mathrm{stoch}}$:
@@
If $h$ is at least comparable to unity (more precisely, $1/\sqrt{2}$; see end of subsubsection), then we consider that broken detailed balance is significant. We may illustrate this criterion in the cases of real or complex eigenvalues of $\mathbf{A}$. For two real eigenvalues $\lambda_{x}$ and $\lambda_{y}$ of $\mathbf{A}$ (Eq. (32)),
@@
For the case of complex conjugate eigenvalues $\lambda\pm i\omega$ of $\mathbf{A}$ with $D^{xy}=0$ (Eq. (34)),
@@ 2.4. Estimation of $\mathbf{A}$ from trajectories
We now consider the estimation of $\mathbf{A}$ in multiple dimensions. It has been shown that for estimation of the force field from stochastic trajectories, there is a finite rate at which information can be extracted [9]. This work has shown that the mean coefficient of the inferred force field $\widehat{\mathbf{A}}\mathbf{x}$ is $\langle\widehat{\mathbf{A}}\rangle=\mathbf{A}$ with correction $\mathcal{O}(T^{-1})$. However, they do not discuss the dependence of $\langle\widehat{\mathbf{A}}\rangle$ on dimension $d$. Eq. (C11) in [9] is summed over basis function index $\alpha$, and we may guess Eq. (C13) in [9] to be of order $\mathcal{O}(d^{3})$, leaving a $\mathcal{O}(d)$ bias for the inferred force field coefficient. To explore the impact of dimension, we explicitly calculate the bias and variance of $\widehat{\mathbf{A}}$ for a $d$-dimensional linear system in which $\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$. According to the information-theoretic criterion in [9], the force field should start to be resolved for a trajectory length of 2 time-units. The value of $\widehat{\mathbf{A}}$ inferred from a trajectory of duration $T$ is [^8] The results in this section hold for inference based on either Itô or Stratonovich calculus in [9]. Also, we have neglected the inference of the true mean $\langle\mathbf{x}\rangle$ and have assumed this as known to be $\mathbf{0}$. This simplifies the calculations while retaining interesting effects.:
@@
The $n=0$ term contributes $\mathds{1}$ to $\langle\widehat{\mathbf{A}}(T)\rangle$. For the contribution of the $n=1$ term, we have from Gaussianity and Isserlis’s theorem (also known as Wick’s theorem [16]):
@@
$\Big.{\qquad{}+\langle\dot{x}^{i}(\tau)x^{l^{\prime}}(\tau^{\prime\prime})\rangle(\langle x^{j}(\tau)x^{k}(\tau^{\prime})\rangle\langle x^{k^{\prime}}(\tau^{\prime\prime})x^{l}(\tau^{\prime})\rangle+\langle x^{j}(\tau)x^{l}(\tau^{\prime})\rangle\langle x^{k}(\tau^{\prime})x^{k^{\prime}}(\tau^{\prime\prime})\rangle)}\Big].$
@@
(52) $\delta^{ij}(\delta^{kk^{\prime}}\delta^{ll^{\prime}}+\delta^{kl^{\prime}}\delta^{lk^{\prime}})\frac{2T+e^{-2T}-1}{2T^{2}}$
@@
Figure 2. Effect of dimensionality on mean of (inferred) $\widehat{A}^{i}_{i}$ (ii not summed) when $\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$.
@@
$\Big.{\qquad{}+\langle\dot{x}^{i}(\tau)x^{l}(\upsilon)\rangle(\langle x^{j}(\tau)\dot{x}^{i^{\prime}}(\tau^{\prime})\rangle\langle x^{j^{\prime}}(\tau^{\prime})x^{k}(\upsilon)\rangle+\langle x^{j}(\tau)x^{j^{\prime}}(\tau^{\prime})\rangle\langle\dot{x}^{i^{\prime}}(\tau^{\prime})x^{k}(\upsilon)\rangle)}\Big].$
@@
$\qquad{}\times(\langle x^{k}(\upsilon)x^{k^{\prime}}(\upsilon^{\prime})\rangle\langle x^{l}(\upsilon)x^{l^{\prime}}(\upsilon^{\prime})\rangle+\langle x^{k}(\upsilon)x^{l^{\prime}}(\upsilon^{\prime})\rangle\langle x^{l}(\upsilon)x^{k^{\prime}}(\upsilon^{\prime})\rangle)+\mathcal{O}\left({\frac{1}{T^{2}}}\right)$
@@
Figure 3. Effect of dimensionality on variance of (inferred) $\widehat{A}^{i}_{i}$ (ii not summed) when $\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$.
@@
Figure 4. Effect of dimensionality on variance of inferred drift ($\widehat{A}^{i}_{k}\widehat{A}^{i}_{l}\widehat{C}^{kl}$, $i$ not summed) when $\mathbf{A}=\mathbf{D}=\mathbf{C}=\mathds{1}$.
@@ 3.1. Preliminaries: Modeling, covariance functions, and detailed balance
 [^10] As with $\mathbf{x}$, we have omitted a possible additive constant in $\dot{y}$. A system with an additive constant may be transformed into Eq. (60) by considering instead the quantity y(t)cty(t)-ct, where $c$ is a constant (in this case, $c=\langle\dot{y}\rangle$). This simplification is again used throughout. Note that such a constant must vanish when detailed balance is satisfied, since we must have in that case $\langle\dot{y}\rangle=0$.
@@ 3.3. Comparison of experimental data with theoretical predictions
where $\Delta_{\textrm{exp\textendash theo}}$ denotes the deviation of experimental to theoretical values. As mentioned for the angular momentum, if the time-step is not small, then on the r.h.s. of the above, it may be suitable to use discrete-time estimators of the diffusion. In case the angular momenta exceed the “reference” values (Eq. (43)) by some factor, we may need to relax the above comparison by the same factor, not just for the antisymmetric part but also for the symmetric part. We may illustrate this by way of two-dimensional linear Gaussian dynamics with complex conjugate eigenvalues $\lambda\pm i\omega$ of $\mathbf{A}$ (Eq. (34)) and $D^{xx}=D^{yy}$, in which case:
@@
Now, we need to extend the notion of $\langle\dot{\mathbf{x}}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$ to integrated variables. We easily make the identification $\text{``$\langle\dot{\mathbf{x}}(\tau)y(0)\rangle$''}=-\langle\mathbf{x}(\tau)\dot{y}(0)\rangle$ for $\tau>0$. (For $\tau=0$, the time-derivative $\dot{y}(0)$ is interpreted in the anti-Itô sense.) For a pair of integrated variables $y$ and $z$ (not the same $z$ as in Eq. (66)), we manipulate the terms in the following equation, similarly to Eq. (8):
(84) $2D^{yz}=\lim_{\tau\to 0^{+}}\frac{\langle(y(\tau)-y(0))(z(\tau)-z(0))\rangle}{\tau}=\text{``$\langle-\dot{y}z-y\dot{z}\rangle$''}.$
Since there is no angular momentum between $y$ and $z$, we assign $\langle\dot{y}z\rangle$ and $\langle y\dot{z}\rangle$ the same (fictional) value, $D^{yz}$. For non-zero time lags, we integrate:
(85) $\text{``$\langle\dot{y}(\tau)z(0)\rangle$''}=\text{``$\langle\dot{y}(0)z(0)\rangle$''}-\int_{0^{+}}^{\tau}\mathrm{d}t\,\langle\dot{y}(t)\dot{z}(0)\rangle.$
The above is equal to $\langle(z(\tau)-z(0))\circ\dot{y}(\tau)\rangle$. From the velocity covariance function $\langle\dot{y}(t)\dot{z}(0)\rangle$ (where $t$ is allowed to run negative), we can compute the expected value of any second-order quantity, such as the two-variable analogue of the mean squared displacement:
@@
where we re-indexed $m=n+k$. Thus, although the estimate Eq. (90) for the velocity autocovariance function is self-averaging for any given finite time lag, its sum over all time lags is not [^11] In the case where $\langle\Delta x\rangle$ is treated as known and can be used in place of $\overline{\Delta x}$ in Eq. (92), the resulting estimate of long-time diffusivity is $(N/2\Delta t)(\overline{\Delta x}-\langle\Delta x\rangle)^{2}$, which has the correct expectation value, but is not self-averaging.. If we denote the limits of summation by $\pm M$, then the limits $M\to\infty$, $N\to\infty$ do not commute. The same considerations apply when integrating the covariance functions of stationary variables.